Evaluate the following integrals. (Show your solution)


"1) \intop 2x√x" "dx"

2) "\intop" "2\/\u221ax\u00b3" "dx"


1
Expert's answer
2022-06-14T00:05:35-0400

We'll use the next basic integration formula:

"\int x^ndx=\frac{x^{n+1}}{n+1}+C."


So we have:

1)

"\int 2x\sqrt{x}dx=\int2x^1x^{\frac{1}{2}}dx=\int2x^{\frac{3}{2}}dx="


"=2\frac{x^{\frac{3}{2}+1}}{\frac{3}{2}+1}+C=2\frac{x^{\frac{5}{2}}}{\frac{5}{2}}+C=\frac{4}{5}x^{{\frac{5}{2}}}+C="


"=\frac{4}{5}x^{2+\frac{1}{2}}+C=\frac{4}{5}x^2x^{\frac{1}{2}}+C=\frac{4}{5}x^2\sqrt{x}+C."


2) "\int\frac{2}{\sqrt{x^3}}dx=\int\frac{2}{x^{\frac{3}{2}}}dx=\int2x^{-\frac{3}{2}}dx="


"=2\frac{x^{-\frac{3}{2}+1}}{-\frac{3}{2}+1}+C=2\frac{x^{-\frac{1}{2}}}{-\frac{1}{2}}+C=-\frac{4}{x^{\frac{1}{2}}}+C=-\frac{4}{\sqrt{x}}+C."


Answer: 1) "\int 2x\sqrt{x}dx=\frac{4}{5}x^2\sqrt{x}+C,"


2) "\int\frac{2}{\sqrt{x^3}}dx=-\frac{4}{\sqrt{x}}+C."